# BadHMP: Backdoor Attack against Human Motion Prediction

### v1

@@ Abstract
Precise future human motion prediction over subsecond horizons from past observations is crucial for various safety-critical applications. To date, only one study has examined the vulnerability of human motion prediction to evasion attacks. In this paper, we propose BadHMP, the first backdoor attack that targets specifically human motion prediction. Our approach involves generating poisoned training samples by embedding a localized backdoor trigger in one arm of the skeleton, causing selected joints to remain relatively still or follow predefined motion in historical time steps. Subsequently, the future sequences are globally modified to the target sequences, and the entire training dataset is traversed to select the most suitable samples for poisoning. Our carefully designed backdoor triggers and targets guarantee the smoothness and naturalness of the poisoned samples, making them stealthy enough to evade detection by the model trainer while keeping the poisoned model unobtrusive in terms of prediction fidelity to untainted sequences. The target sequences can be successfully activated by the designed input sequences even with a low poisoned sample injection ratio. Experimental results on two datasets (Human3.6M and CMU-Mocap) and two network architectures (LTD and HRI) demonstrate the high-fidelity, effectiveness, and stealthiness of BadHMP. Robustness of our attack against fine-tuning defense is also verified.
@@ I INTRODUCTION
Human motion prediction is a sequence-to-sequence task where future motion sequences are predicted based on observed historical motion sequences. Accurate forecasting of future human poses is crucial for the success of various applications, such as human-robot interaction and collaboration (HRI/C) [1, 2], human tracking [3], autonomous driving [4], and rehabilitation [5].
Various advanced neural network architectures have been explored for this task, including recurrent neural networks (RNNs) [6, 7, 8], graph convolutional networks (GCNs) [9, 10, 11], generative models [12, 13, 14, 15], and Transformers [16, 17]. Despite the extensive research into deep learning based human motion prediction, the vulnerability of these models to potential attacks has not been sufficiently explored. To date, only one evasion attack [18] has been investigated on convolution-based human motion prediction. Hence, there exists a significant gap in understanding the robustness of human motion prediction models against other forms of malicious attacks.
Backdoor attack is an emerging form of data-poisoning attack targeting deep neural networks (DNNs), where the adversary subtly alters a small subset of training samples by embedding a trigger into the input data and substituting the corresponding outputs with predefined targets. During training, the victim model inadvertently learns both the intended tasks and a strong association between the trigger and the target output. At the inference stage, the model behaves as a benign model under normal conditions but consistently produces the predefined target output when the trigger is present in the input.
Most existing backdoor attacks focus on image classification tasks [19, 20, 21, 22, 23], while a few have been extended to other tasks [24, 25, 26, 27, 28]. However, backdoor attacks on human motion prediction models remain uncharted. A successful backdoor attack in this domain poses safety hazards that may lead to grave consequences. For instance, in the scenario of HRI/C, a robot equipped with a poisoned model may inaccurately predict human motions, leading to erroneous decisions and potentially hazardous outcomes in subsequent time steps. The main challenges of launching a backdoor attack on human motion prediction are as follows: 1) Due to the unique data format of human motion samples (spatial and temporal 3D joint positions), existing data-poisoning techniques are not directly applicable for generating the poisoned samples for such task; 2) To avoid detection by the model trainer, the poisoned training samples of human motion sequences must remain smooth and natural. This means that the clean samples need to be subtly manipulated to ensure that the fundamental physics principles of human body dynamics are not violated.
@@
  • To the best of our knowledge, this is the first work that investigates the vulnerability of human motion prediction models to backdoor attacks;
  • Two new types of backdoor triggers (“rigid” and “soft”) and two types of targets (“jammed” and “moving”) dedicated to human motion prediction are designed, resulting in four distinct poisoning strategies;
  • Two novel evaluation metrics, Clean Data Error (CDE) and Backdoor Data Error (BDE), are proposed to assess the attack performance on human motion predictors;
  • Extensive experiments are conducted on two popular benchmark datasets (Human3.6M and CMU-Mocap) and two widely used model architectures (LTD and HRI) to attest the robustness of our proposed attack.
@@ II-A Human Motion Prediction
The GCN-based motion prediction method was first introduced in  [9] by employing the Discrete Cosine Transform (DCT) to encode the spatial dependency and temporal information of human poses. This approach was further refined [10] by capturing similarities between current and historical motion contexts. Inspired by the success of GCN in modeling dynamic relations among pose joints [9], various GCN-based prediction methods [11, 33] have been developed for more complex spatio-temporal dependencies over diverse action types. Recently, the attention mechanism [16, 17] of Transformers [34] has been leveraged to capture the spatial, temporal, and pairwise joint relationships within motion sequences. Without resorting to complex deep learning architectures, a lightweight multi-layer perception combined with DCT and standard optimization techniques can achieve excellent performance with fewer parameters [35]. Sampling from deep generative models [12, 13, 14, 15] trained over large motion-capture dataset has also been devised for more realistic and coherent stochastic human motion prediction.
@@ II-B Backdoor Attacks and Defenses
Backdoor attacks train the victim model with poisoned training data to embed a malicious backdoor that can be activated at test time to cause the victim model to misbehave. The most popular attack dates back to BadNets [19], where a small number of training samples are stamped with a tiny fixed binary pattern (a.k.a trigger) at the right bottom corner and relabeled to a target class. The backdoor feature is learnt by the DNN classifier during the training process. Since then, a series of improved backdoor attacks have been developed, employing techniques such as image blending transformations [36], steganography [21], warping transformations [20], and adaptive optimizations [37], to enhance the stealthiness of the backdoor trigger, thereby evading detection by the model trainer. An even more inconspicuous branch of clean-label backdoor attacks [23, 22, 38] can achieve target misclassification by hiding the backdoor triggers into the training images without altering their labels. While most existing backdoor attacks have been developed for image classification tasks, some successful attacks have also been reported in other task domains, such as speech recognition [24, 25], graph classification [26, 27], crowd counting [28], and point cloud classification [39, 40].
@@ III Threat Model
Given a history motion sequence $\textbf{X}_{1:N}=\left[\textbf{x}_{1},\textbf{x}_{2},\cdots,\textbf{x}_{N}\right]$ that is composed of $N$ consecutive frames of human poses, the human motion prediction model $f_{\theta}$ parameterized by $\theta$ aims to forecast the future $T$ frames of poses as $\textbf{X}_{N+1:N+T}=\left[\textbf{x}_{N+1},\textbf{x}_{N+2},\cdots,\textbf{x}_{N+T}\right]$, where each pose $\textbf{x}_{i}\in\mathbb{R}^{K\times 3}$ consists of 3D coordinates of $K$ joints. Let $D_{\text{tr}}$ and ${D}_{\text{ts}}$ denote the training and test datasets, respectively. The training process aims to solve the following optimization problem:
$\theta^{\ast}$ $=\mathop{\arg\min}\mathbb{E}_{\textbf{X}\sim D_{\text{tr}}}\left[\pazocal{L}_{m}(\hat{\textbf{X}}_{N+1:N+T},\textbf{X}_{N+1:N+T})\right],$ (1)
$=\mathop{\arg\min}\mathbb{E}_{\textbf{X}\sim D_{\text{tr}}}\left[\frac{1}{K\times T}\sum_{n=1}^{T}\sum_{j=1}^{K}\left\|\hat{\textbf{x}}_{(j,n)}-\textbf{x}_{(j,n)}\right\|^{2}\right],$
where $\hat{\textbf{X}}_{N+1:N+T}=f_{\theta}(\textbf{X}_{1:N})$ and $\textbf{X}_{N+1:N+T}$ denote the predicted and ground-truth poses of future $T$ time steps, respectively. $\hat{\textbf{x}}_{(j,n)}$ represents the predicted $j$-th joint position at frame $n$, and $\textbf{x}_{(j,n)}$ is its corresponding ground turth. $\pazocal{L}_{m}$ denotes the Mean Per Joint Position Error (MPJPE) [48] in millimeter, which is the most widely used metric for 3D pose error evaluation.
@@ III-A Attack Scenario
Following the threat model of backdoor attacks on image classification models  [19, 20, 23], we assume that the attacker is a malicious third party who provides the training set to the model trainer. In this scenario, the attacker is allowed to poison $\rho$of samples of the clean training set $D_{\text{tr}}$ before the training stage, where $\rho=(N_{\text{poison}}/N_{\text{train}})\times 100\%$ is commonly referred to as the injection ratio. However, the attacker has no knowledge of or access to the training pipeline, including the model architecture, optimization algorithm, training loss, etc. Additionally, manipulating the well-trained victim model is also not permitted. The backdoor sample generation process can be expressed as $=G(\textbf{X})$ X ), where $G(\cdot)$ denotes the poisoning function that will be elaborated in Sec. IV, and $\tilde{\textbf{X}}$ is the poisoned sample.
@@ III-B Attacker’s Goals
$\text{CDE}(f)=\mathbb{E}_{\textbf{X}\sim D_{\text{ts}}}\left[\pazocal{L}_{m}(f(\textbf{X}_{1:N}),\textbf{X}_{N+1:N+T})\right].$ (2)
The CDE of the victim model should be comparable to that of the benign model, i.e., $\lvert\text{CDE}(f_{\theta^{\prime}})-\text{CDE}(f_{\theta})\rvert\leq\epsilon$, with $\epsilon$ being a samll positive threshold.
@@
$\text{BDE}(f)=\mathbb{E}_{\tilde{\textbf{X}}\sim\tilde{D}_{\text{ts}}}\left[\pazocal{L}_{m}(f(\tilde{\textbf{X}}_{1:N}),\tilde{\textbf{X}}_{N+1:N+T})\right],$ (3)
@@
Stealthiness. Poisoned samples should look similar to its clean version to avoid being detected by human inspector or automatic checker. Specifically, the poisoned pose sequences should be smooth and natural. Mean per-joint acceleration (Acc) and jerk (second- and third-order derivatives of the joint positions, respectively) for human motion synthesis [49, 50] are utilized to evaluate the smoothness of the poisoned samples. Moreover, inspired by the physics-constrained attack [18], we compute the change of bone length to evaluate the naturalness. There exists a bone between a pair of connected joints, and the bone length change is small during the motion as human bones are not elastic. Hence, the bone length change (BLC) of poisoned samples should be kept as low as that of their clean versions. The above three metrics are formulated as follows:
@@
where $S_{l,n}$ denotes the length of the $l$-th bone at frame $n$, and $L_{C}$ is the total number of bones. $\ddot{\textbf{X}}$ and $\overset{\text{...}}{\textbf{X}}$ denote the second- and third-order derivatives of X, respectively.
Figure 1: Paired clean (solid) and poisoned (dotted) training samples. The poisoning strategies are R-J (Row 1), R-M (Row 2), S-J (Row 3), and S-M (Row 4), respectively. The clean training samples are randomly selected from the action “walking” of the CMU dataset, and the trajectories leveraged for poisoning is extracted from a randomly selected test sample of the “basketball” action. Due to the page limit, the complete 75-frame motion sample is down-sampled to 18 frames for display in this figure, with 10 frames for input and 8 frames for output.
@@ IV The Proposed Attack
Our attack consists of three steps: localized history sequences poisoning, global future sequences poisoning, and sample selection.
@@ IV-A Localized History Sequences Poisoning
A body pose is composed of five parts: torso, left arm, right arm, left leg, and right leg. On the $N$-frame input sequences, only several connected joints in a fixed limb are manipulated for backdoor trigger embedding, while the remaining joints’ positions are unchanged. This is to ensure that the semantic meaning of the input sequences is not damaged. For instance, a “walking” sample can be poisoned by controlling the “left arm” to do a predefined motion like staying rigid or drawing a circle. In this paper, we select $m$ joints on the left arm for poisoning, denoted as $J_{p}=\{j_{1},j_{2},\cdots,j_{m}\}$. The joint of the left shoulder is $j_{0}$.
Two types of triggers are proposed: rigid and soft.
Rigid trigger creates a condition where the joints on one arm (the non-dominant left arm is chosen for convenience) form a predefined pose and remain relatively still in the 3D space, with movements closely tied to $j_{0}$. Hence, the left arm appears “frozen” during the input time steps. Given a source sample $\textbf{X}^{\text{src}}$ randomly selected from the clean test set, the left arm pose at the first frame is grafted onto the $N$ input frames of the clean training sample X. However, instead of directly grafting the left arm, a scaling transformation is first applied to $\textbf{X}^{\text{src}}$, producing $^{\text{src}}=\text{Scale}(\textbf{X}^{\text{src}},\textbf{X})$ X ), where the bone lengths match those in X but the directions remain the same as in $\textbf{X}^{\text{src}}$. This step avoids the disproportionate grafted left arm. The grafting operation is then performed as:
$\tilde{\textbf{x}}_{(j,n)}=\begin{cases}\textbf{x}_{(j_{0},n)}+\left[\overline{\textbf{x}}^{\text{src}}_{(j,1)}-\overline{\textbf{x}}^{\text{src}}_{(j_{0},1)}\right],\qquad&\text{if}\ j\in J_{p},\\ \textbf{x}_{(j,n)},&\text{otherwise},\end{cases}$ (7)
where $1\leq n\leq N$. The rigid trigger embedding process can be regarded as adopting a predefined left arm pose to the training input sequences without changing the bone lengths.
Soft trigger also grafts the left arm pose from the source sample to the training sample. However, unlike the rigid trigger, it involves a frame-by-frame basis left arm pose grafting throughout the input time steps. Consequently, the left arm in the poisoned sample exhibits a predefined motion naturally and flexibly. This operation can be formulated as:
$\tilde{\textbf{x}}_{(j,n)}=\begin{cases}\textbf{x}_{(j_{0},n)}+\left[\overline{\textbf{x}}^{\text{src}}_{(j,n)}-\overline{\textbf{x}}^{\text{src}}_{(j_{0},n)}\right],\qquad&\text{if}\ j\in J_{p},\\ \textbf{x}_{(j,n)},&\text{otherwise}.\end{cases}$ (8)
Since the left arm motion is extracted from a real test sample, its smoothness and naturalness are guaranteed. Generally, the soft trigger is stealthier than rigid trigger because the manipulated left arm continues to move flexibly, thus is less likely to draw attention. However, poisoning the victim model with the rigid trigger makes it easier for the attacker to activate the embedded backdoor at the inference stage - the attacker simply needs to ensure that the left arm maintains a predefined pose for $N$ frames to induce the incorrect prediction.

### v2

@@ Abstract
Precise future human motion prediction over sub-second horizons from past observations is crucial for various safety-critical applications. To date, only a few studies have examined the vulnerability of skeleton-based neural networks to evasion and backdoor attacks. In this paper, we propose BadHMP, a novel backdoor attack that targets specifically human motion prediction tasks. Our approach involves generating poisoned training samples by embedding a localized backdoor trigger in one limb of the skeleton, causing selected joints to follow predefined motion in historical time steps. Subsequently, the future sequences are globally modified that all the joints move following the target trajectories. Our carefully designed backdoor triggers and targets guarantee the smoothness and naturalness of the poisoned samples, making them stealthy enough to evade detection by the model trainer while keeping the poisoned model unobtrusive in terms of prediction fidelity to untainted sequences. The target sequences can be successfully activated by the designed input sequences even with a low poisoned sample injection ratio. Experimental results on two datasets (Human3.6M and CMU-Mocap) and two network architectures (LTD and HRI) demonstrate the high-fidelity, effectiveness, and stealthiness of BadHMP. Robustness of our attack against fine-tuning defense is also verified.
@@ I INTRODUCTION
Human motion prediction is a sequence-to-sequence task where future motion sequences are predicted based on observed historical motion sequences. Accurate forecasting of future human poses is crucial for the success of various applications, such as human-robot interaction and collaboration (HRI/C) [1, 2], human tracking [3], autonomous driving [4], and particularly in healthcare and biomedical fields, such as seamless interactions with exoskeletons and prosthetic devices that enable more effective rehabilitation [5].
Various advanced neural network architectures have been explored for this task, including recurrent neural networks (RNNs) [6, 7, 8], graph convolutional networks (GCNs) [9, 10, 11], generative models [12, 13, 14, 15], and Transformers [16, 17]. Despite the extensive research into deep learning based human motion prediction, the vulnerability of these models to potential attacks has not been sufficiently explored. To date, only a few evasion attacks [18, 19] have been investigated on human motion prediction. Hence, there exists a significant gap in understanding the robustness of human motion prediction models against other forms of malicious attacks.
Backdoor attack targeting deep neural networks (DNNs) is a form of data-poisoning attack, where the adversary subtly alters a small subset of training samples by embedding a trigger into the input data and substituting the corresponding outputs with predefined targets. During training, the victim model inadvertently learns both the intended tasks and a strong association between the trigger and the target output. At the inference stage, the model behaves as a benign model under normal conditions but consistently produces the predefined target output when the trigger is present in the input.
Most existing backdoor attacks focus on image classification tasks [20, 21, 22, 23, 24], while some studies have been extended to other tasks [25, 26, 27, 28, 29]. While backdoor attacks have been studied in other skeleton-data-based machine leaning tasks [30, 31], they have not been explored in the context of human motion prediction models. A successful backdoor attack in this domain poses safety hazards that may lead to grave consequences. For instance, in the scenario of HRI/C, a robot equipped with a poisoned model may inaccurately predict human motions, leading to erroneous decisions and potentially hazardous outcomes in subsequent time steps. The main challenges of launching a backdoor attack on human motion prediction are as follows: 1) Due to the unique data format of human motion samples (spatial and temporal 3D joint positions), existing data-poisoning techniques are not directly applicable for generating the poisoned samples for such task; 2) To avoid detection by the model trainer, the poisoned training samples of human motion sequences must remain smooth and natural. This means that the clean samples need to be subtly manipulated to ensure that the fundamental physics principles of human body dynamics are not violated.
@@
  • We propose a novel poisoning strategy that generates smooth and natural adversarial human motion samples. Specifically, we extract the motion of a selected limb from a source sample, and graft this predefined motion onto clean samples to seamlessly embed the backdoor trigger into input sequences. For output sequences, we globally extract and transfer the trajectories of all joints from the source sample to clean samples as the target motion patterns.
  • We design two novel evaluation metrics, Clean Data Error (CDE) and Backdoor Data Error (BDE), to assess the attack performance on human motion predictors.
  • Extensive experiments are conducted on two popular benchmark datasets (Human3.6M and CMU-Mocap) and two widely used model architectures (LTD and HRI) to attest the performance of our proposed attack.
@@ II-A Human Motion Prediction
The GCN-based motion prediction method was first introduced in [9] by employing the Discrete Cosine Transform (DCT) to encode the spatial dependency and temporal information of human poses. This approach was further refined by capturing similarities between current and historical motion contexts [10]. Inspired by the success of GCN in modeling dynamic relations among pose joints [9], various GCN-based prediction methods [11, 36] have been developed for more complex spatio-temporal dependencies over diverse action types. Recently, the attention mechanism [16, 17] of Transformers [37] has been leveraged to capture the spatial, temporal, and pairwise joint relationships within motion sequences. Without resorting to complex deep learning architectures, a lightweight multi-layer perception combined with DCT and standard optimization techniques can achieve excellent performance with fewer parameters [38]. Sampling from deep generative models [12, 13, 14, 15] trained over large motion-capture dataset has also been devised for more realistic and coherent stochastic human motion prediction.
@@ II-B Backdoor Attacks and Defenses
Backdoor attacks train the victim model with poisoned training data to embed a malicious backdoor that can be activated at test time to cause the victim model to misbehave. The most popular attack dates back to BadNets [20], where a small number of training samples are stamped with a tiny fixed binary pattern (a.k.a trigger) at the right bottom corner and relabeled to a target class. The backdoor feature is learnt by the DNN classifier during the training process. Since then, a series of improved backdoor attacks have been developed, employing techniques such as image blending transformations [39], steganography [22], warping transformations [21], and adaptive optimizations [40], to enhance the stealthiness of the backdoor trigger, thereby evading detection by the model trainer. An even more inconspicuous branch of clean-label backdoor attacks [24, 23, 41] can achieve target misclassification by hiding the backdoor triggers into the training images without altering their labels. While most existing backdoor attacks have been developed for image classification tasks, some successful attacks have also been reported in other task domains, such as speech recognition [25, 26], graph classification [27, 28], skeleton action recognition [31], and human pose estimation [30] .
@@ III Threat Model
Given a history motion sequence $X_{1:N}=\left[X_{1},X_{2},\cdots,X_{N}\right]$ that is composed of $N$ consecutive frames of human poses, the human motion prediction model $f_{\theta}$ parameterized by $\theta$ aims to forecast the future $T$ frames of poses as $X_{N+1:N+T}=\left[X_{N+1},X_{N+2},\cdots,X_{N+T}\right]$, where each pose $X_{i}\in\mathbb{R}^{K\times 3}$ consists of 3D coordinates of $K$ joints. Let $D_{\text{tr}}$ and ${D}_{\text{ts}}$ denote the training and test datasets, respectively. The training process aims to solve the following optimization problem:
$\theta^{\ast}$ $=\mathop{\arg\min}\mathbb{E}_{X\sim D_{\text{tr}}}\left[\pazocal{L}_{m}(\hat{X}_{N+1:N+T},X_{N+1:N+T})\right],$ (1)
$=\mathop{\arg\min}\mathbb{E}_{X\sim D_{\text{tr}}}\left[\frac{1}{K\times T}\sum_{n=1}^{T}\sum_{j=1}^{K}\left\|\hat{X}_{(j,n)}-X_{(j,n)}\right\|^{2}\right],$
where $\hat{X}_{N+1:N+T}=f_{\theta}(X_{1:N})$ and $X_{N+1:N+T}$ denote the predicted and ground-truth poses of future $T$ time steps, respectively. $\hat{X}_{(j,n)}$ represents the predicted $j$-th joint position at frame $n$, and $X_{(j,n)}$ is its corresponding ground truth. $\pazocal{L}_{m}$ denotes the Mean Per Joint Position Error (MPJPE) [49] in millimeter, which is the most widely used metric for 3D pose error evaluation.
@@ III-A Attack Scenario
Following the threat model of backdoor attacks on image classification models [20, 21, 24], we assume that the attacker is a malicious third party who provides the training set to the model trainer. In this scenario, the attacker is allowed to poison $\rho$of samples of the clean training set $D_{\text{tr}}$ before the training stage, where $\rho=(N_{\text{poison}}/N_{\text{train}})\times 100\%$ is commonly referred to as the injection ratio. However, the attacker has no knowledge of or access to the training pipeline, including the model architecture, optimization algorithm, training loss, etc. Additionally, manipulating the well-trained victim model is also not permitted. The backdoor sample generation process can be expressed as $\tilde{X}=G(X)$, where $G(\cdot)$ denotes the poisoning function that will be elaborated in Sec. IV, and $\tilde{X}$ is the poisoned sample.
@@ III-B Attacker’s Goals
$\text{CDE}(f)=\mathbb{E}_{X\sim D_{\text{ts}}}\left[\pazocal{L}_{m}(f(X_{1:N}),X_{N+1:N+T})\right].$ (2)
The CDE of the victim model should be comparable to that of the benign model, i.e., $\lvert\text{CDE}(f_{\theta^{\prime}})-\text{CDE}(f_{\theta})\rvert\leq\epsilon$, with $\epsilon$ being a small positive threshold.
@@
$\text{BDE}(f)=\mathbb{E}_{\tilde{X}\sim\tilde{D}_{\text{ts}}}\left[\pazocal{L}_{m}(f(\tilde{X}_{1:N}),\tilde{X}_{N+1:N+T})\right],$ (3)
@@
Figure 1: Row 1: the source sample with the semantic meaning of “walking”. Only joints in green are leveraged to generate the trigger or target. Row 2: a clean sample with the semantic meaning of “soccer”. Row 3: the poisoned version of the above clean sample. Row 4: Comparison of the paired clean (solid) and poisoned (dotted) samples. Due to the page limit, the complete 75-frame motion sample is down-sampled to 15 frames for display in this figure, with 10 frames for input and 5 frames for output.
Stealthiness. Poisoned samples should look similar to its clean version to avoid being detected by human inspector or automatic checker. Specifically, the poisoned pose sequences should be smooth and natural. Mean per-joint acceleration (Acc) and jerk (second- and third-order derivatives of the joint positions, respectively) for human motion synthesis [50, 51] are utilized to evaluate the smoothness of the poisoned samples. Moreover, following the physics-constrained attack [18], we compute the change of bone length to evaluate the naturalness. There exists a bone between a pair of connected joints (e.g., the humerus bone between shoulder and elbow), and the bone length change is small during the motion as human bones are not elastic. Hence, the bone length change (BLC) of poisoned samples should be kept as low as that of their clean versions. The above three metrics are formulated as follows:
@@
where $S_{l,n}$ denotes the length of the $l$-th bone at frame $n$, and $L_{C}$ is the total number of bones.
@@ IV The Proposed Attack
Our attack consists of three stages: (1) localized history sequences poisoning, (2) global future sequences poisoning, and (3) victim model poisoning.
@@ IV-A Localized History Sequences Poisoning
A body pose is composed of five parts: torso, left arm, right arm, left leg, and right leg. On the $N$-frame input sequences, only several connected joints in a selected limb are manipulated for backdoor trigger embedding, while the remaining joints’ positions are unchanged. This is to ensure that the semantic meaning of the input sequences is not damaged.
By default, we select the left arm for poisoning. Given a clean sample $X$ to be poisoned, and a source sample $X^{\text{src}}$ randomly selected from the clean test set. For the $N$ input frames, we graft the left arm pose of $X^{\text{src}}$ onto that of $X$. For instance, a “walking” sample can be poisoned by controlling the left arm to do a predefined motion grafted from the source sample like “drawing a circle”. However, since the clean and source samples may captured from actors of different physiques, direct grafting may causing disproportionate grafted left arm. Therefore, a 3D scaling transformation is first applied to $X^{\text{src}}$, producing $\overline{X}^{\text{src}}=\text{Scale}(X^{\text{src}},X)$, where the bone lengths match those in $X$ but the directions remain the same as in $X^{\text{src}}$.
The left arm to be poisoned is formed by $m$ joints denoted as $J_{p}=\{j_{1},j_{2},\cdots,j_{m}\}$, and the joint of the left shoulder is $j_{0}$. The grafting operation is performed as:
$\tilde{X}_{(j,n)}=\begin{cases}X_{(j_{0},n)}+\left[\overline{X}^{\text{src}}_{(j,n)}-\overline{X}^{\text{src}}_{(j_{0},n)}\right],\qquad&\text{if}\ j\in J_{p},\\ X_{(j,n)},&\text{otherwise},\end{cases}$ (7)
where $1\leq n\leq N$. This trigger embedding process can be regarded as adopting a predefined left arm motion to the clean training input sequences without changing the bone lengths. Since the left arm motion is extracted from a real source sample, its smoothness and naturalness are guaranteed.